### v2

@@ B.1 Action Space
  • Divide and Conquer ($a_{5}$): Breaking down a complex reasoning problem into several smaller subproblems and progressively solving them to achieve the overall solution.
  • Self-Reflection ($a_{6}$): Engaging in timely reflection of prior solutions and implementing necessary refinement during the reasoning process to ensure accuracy.
Table 8: Comparison with other search-based LLM and MLLM methods. Note that, most method contain limited space. In contrast, we define a rich set of reasoning actions, thus enhancing the upper bound of model reasoning capabilities.
@@
Visual Question Answering ChartQA Masry et al. (2022) chart understanding and reasoning
MMStar Chen et al. (2024b) 6 core capabilities, like scientific reasoning
Mathematics MathVista Lu et al. (2023) 12 core capabilities, like arithmetic reasoning
MathVerse Zhang et al. (2025) 6 distinct versions-text-dominant
MathVision Wang et al. (2024a) 16 mathematical domains
Commonsense and science GAOKAO-MM Zong & Qiu (2024) 8 subjects, like history
B.2 Reward Value in MCTS
To simplify the construction of thought cards via the MCTS iteration, we avoid introducing an external reward model to score each step. Given the current skepticism regarding the self-rewarding capabilities of language models, alternative methods are necessary. Inspired by the principle that actions leading to correct answers should be rewarded more frequently, we aim to increase their likelihood of selection in future MCTS tree expansions. Following Zhou et al. (2024); Wu et al. (2024a), we define the reward value as the likelihood (or confidence) of self-consistency via majority voting. Note that this principle applies only to leaf nodes. The Q-values of intermediate nodes are initially set to 0 and are dynamically updated during the backpropagation process, as described in Equation 1.
Appendix C More Details about Experimental Setup
C.1 Benchmarks and Datasets
@@
  • ChartQA Masry et al. (2022) is a comprehensive benchmark for chart-based question answering that emphasizes visual and logical reasoning capabilities. The dataset comprises 9,608 human-authored questions and 23,111 machine-generated questions derived from human-written chart summaries. It incorporates 20,882 real-world charts from diverse sources including Statista, Pew Research, Our World in Data, and OECD, spanning multiple domains and visualization styles. ChartQA addresses prior dataset limitations through its focus on human-authored questions, real-world charts, and open-vocabulary responses. We utilize this dataset to evaluate our method’s performance on visual chart question-answering tasks commonly encountered in real-world scenarios.
@@
  • MMStar Chen et al. (2024b) serves as a vision-critical multimodal benchmark for evaluating MLLM capabilities. It addresses two fundamental limitations in existing evaluations: the redundancy of visual inputs and potential training data contamination. Comprising 1,500 meticulously curated challenge samples across six core competencies and 18 evaluation axes, the benchmark undergoes rigorous validation to ensure visual necessity and minimal data leakage. We employ this comprehensive dataset to assess our method’s general reasoning capabilities.
@@
  • MathVista Lu et al. (2023) is a mathematical visual benchmark consisting of 6,141 examples. These examples are divided into two subsets: testmini with 1,000 annotated examples and test comprising 5,141 problems without public answers. We conduct our evaluation on the testmini subset to assess visual comprehension and compositional reasoning capabilities.
@@
  • MathVerse Zhang et al. (2025) is a comprehensive and specialized visual mathematics benchmark for assessing multimodal mathematical reasoning capabilities of MLLMs. The benchmark contains 2,612 visual math problems, combining 1,236 newly collected problems from public repositories with 1,376 problems from existing benchmarks. Human annotators have transformed each problem into six distinct versions-text-dominant, text-lite, text-only, vision-intensive, vision-dominant, and vision-only—each offering different levels of multimodal information. We utilize this dataset to evaluate AStar’s performance across varying degrees of multimodal integration.
@@
  • MathVision Wang et al. (2024a) constitutes a meticulously compiled collection of 3,040 mathematics problems, each incorporating visual elements from authentic mathematics competitions. The dataset spans 16 mathematical domains and features a five-tier difficulty classification system. We employ this dataset to evaluate our model’s performance on challenging reasoning tasks.
@@
  • GAOKAO-MM Zong & Qiu (2024) represents a multimodal evaluation framework derived from the Chinese National College Entrance Examination (Gaokao). It encompasses eight academic subjects and includes twelve categories of images, such as diagrams, function graphs, maps, and photographs. The benchmark aims to evaluate models’ abilities to understand and reason over diverse multimodal content, reflecting the complexity and breadth of knowledge. We leverage this dataset to assess AStar’s cross-domain generalization capabilities.
C.2 Baselines
@@
Open-source general MLLMs: Recent advances include Qwen2-VL Wang et al. (2024b) and InternVL2 Chen et al. (2024d), which demonstrate exceptional capabilities in complex reasoning and mathematical problem-solving.
@@
      G-LLaVA Gao et al. (2023): Introduces the Geo170K dataset, comprising over 170K geometric image-caption and question-answer pairs, to enhance MLLMs’ geometric problem-solving capabilities.
@@
      Math-LLaVA Shi et al. (2024): Develops the MathV360K dataset by collecting 40K high-quality images with question-answer pairs and synthesizing an additional 320K pairs, then fine-tunes the LLaVA-1.5 model to improve multimodal mathematical reasoning.
@@
      MultiMath Peng et al. (2024): Constructs the MultiMath-300K dataset, encompassing K-12 level mathematical problems with image captions and step-wise solutions, and trains the MultiMath-7B model to bridge visual and mathematical reasoning.
@@
      Math-PUMA Zhuang et al. (2024): Proposes a three-stage Progressive Upward Multimodal Alignment methodology to enhance MLLMs’ mathematical reasoning skills, focusing on aligning visual and textual modalities through a structured training process.
@@
      LlamaV-o1 Thawakar et al. (2025): Introduces a multimodal visual reasoning model trained using a multi-step curriculum learning approach, organizing tasks progressively to facilitate incremental skill acquisition and problem-solving.
@@
      LLaVA-CoT Xu et al. (2024): Develops a vision-language model designed to conduct autonomous multistage reasoning, employing a structured approach that includes summarization, visual interpretation, logical reasoning, and conclusion generation.
@@
      URSA Luo et al. (2025): Proposes a three-module synthesis strategy integrating CoT distillation, trajectory-format rewriting, and format unification, resulting in the MMathCoT-1M dataset and the URSA-7B model, which demonstrates state-of-the-art performance on multiple multimodal mathematical benchmarks.
@@
      InfiMM-Math Han et al. (2024): Introduces InfiMM-WebMath-40B, a high-quality dataset comprising 24 million web pages, 85 million associated image URLs, and 40 billion text tokens, aiming to improve MLLMs’ mathematical reasoning through large-scale multimodal pre-training.
@@
    Slow-Thinking System Implementation:
@@
      Virgo Du et al. (2025): Explores a straightforward approach to implementing multimodal slow-thinking systems by fine-tuning a capable MLLM with a small amount of textual long-form thought data, demonstrating that such reasoning processes can be effectively transferred to MLLMs.
Advanced closed-source MLLMs: Leading proprietary models including GPT-4V OpenAI (2023), GPT-4o OpenAI (2024), and Gemini-1.5-Pro Team et al. (2023) demonstrate exceptional capabilities in multimodal understanding and task-solving, setting benchmarks for open-source alternatives.
Multimodal tree-based methods: Recent works incorporate explicit search mechanisms into multimodal reasoning. AR-MCTS Dong et al. (2024a) enhances reasoning by combining Monte Carlo Tree Search (MCTS) with active retrieval, improving reasoning space diversity and reliability. Mulberry Yao et al. (2024) leverages multi-model collaboration through MCTS’s four iterative operations (selection, expansion, simulation, backpropagation) to identify optimal reasoning paths.
C.3 Implementation Details
@@
In the MCTS-powered prior thought card construction stage, we implement an early termination strategy based on self-consistency Wang et al. (2023) for enhanced efficiency. Building on the observation that repeated action sampling at the same state often indicates successful task completion, we terminate simulation early when the model’s consistency score exceeds a threshold $c$ (i.e., $SC(s)>c$).
@@
Appendix D Supplementary Results
This section presents supplementary results and analysis, including: D.1 Detailed Results on Multimodal Reasoning Benchmarks, D.2 Comparison with Strong Baselines, D.3 Integration with SFT, and D.4 Weak-to-Strong Generalization.
D.1 Detailed Results on Multimodal Reasoning Benchmarks
@@
mPLUG-Owl2-7B Ye et al. (2023) 7B 22.2 23.6 19.2 23.9 13.5 12.7 26.3 21.4
LLaVA-1.5-13B Liu et al. (2024b) 13B 25.7 19.6 28.6 17.6 10.8 27.8 33.6 22.9
MiniGPT-v2 Chen et al. (2023) 7B 23.1 28.1 21.0 24.7 16.2 16.7 25.4 17.9
InternLM-XComposer2-VL-7B Dong et al. (2024c) 7B 47.8 32.0 51.6 30.5 13.5 43.8 37.7 62.8
SPHINX-MoE Lin et al. (2023) 8$\times$ 7B 42.3 31.7 41.6 30.5 16.2 27.1 50.8 50.8
DeepSeek-VL Lu et al. (2024) 7B 34.9 29.2 38.8 27.2 18.9 43.1 35.3 33.2
InternVL2-8B Chen et al. (2024d) 8B 58.3 59.8 56.4 60.3 10.8 30.6 59.0 68.8
Qwen2-VL Wang et al. (2024b) 7B 58.9 44.1 57.5 43.1 24.3 41.7 66.4 75.1
@@
G-LLaVA Gao et al. (2024) 7B 25.1 36.0 19.4 37.6 15.2 17.7 21.0 15.1
Math-LLaVA Shi et al. (2024) 13B 46.6 51.5 40.7 56.2 23.3 34.7 47.7 42.3
Multimath-7B Peng et al. (2024) 7B 50.0 61.9 42.2 64.9 23.3 32.6 42.6 49.2
Math-PUMA-Qwen2-7B Zhuang et al. (2024) 7B 47.9 47.7 46.2 47.3 21.6 32.6 42.6 55.8
URSA-7B Luo et al. (2025) 7B 59.8 74.0 53.5 77.4 21.6 35.4 58.2 57.1
AStar (Ours, Training-free Reasoning) 7B 63.5 69.0 63.1 71.7 61.3 60.0 48.2 68.8
@@
D.2 Comparison with Strong Baselines
@@
We also benchmark our approach against recent powerful multimodal reasoning methods, including LLaVA-CoT Xu et al. (2024) and Virgo Du et al. (2025). As shown in Table 12, comparative experiments across different model scales demonstrate that our AStar reasoning framework, which effectively integrates MLLMs’ inherent reasoning capabilities with external reasoning guidelines, achieves superior performance with minimal prior data. On challenging datasets like MathVerse, our method exhibits substantial improvements, with the 2B model achieving comparable performance to the extensively trained Virgo 7B model. Further analysis reveals two key insights: First, methods relying on vast solution spaces often struggle to identify appropriate reasoning paths, analogous to finding a needle in a haystack. Second, approaches dependent on large-scale training data typically face difficulties in fully capturing complex long-chain reasoning patterns through implicit learning. In contrast, our method adaptively identifies suitable reasoning strategies based on problem complexity, enabling efficient inference across diverse scenarios.
@@
GPT-4o-0513 OpenAI (2024) 63.8 50.2 30.4 48.2
GPT-4V OpenAI (2023) 49.9 54.4 24.8 43.1
Gemini-1.5-Pro Team et al. (2023) 63.9 35.3 19.2 39.5
Claude-3.5-Sonnet Anthropic (2024) 67.7 - - -
Qwen-VL-Plus Bai et al. (2023) 43.3 21.3 10.8 25.2
@@
Qwen2-VL-72B Wang et al. (2024b) 70.5 - 25.9 -
LLaVA-OneVision-72B Li et al. (2024) 67.5 39.1 - -
InternVL2.5-26B Chen et al. (2024d) 67.7 40.1 28.0 45.3
InternVL2-Llama3-76B Chen et al. (2024d) 65.5 42.8 23.7 44.0
Astar (Ours) 63.5 54.0 32.4 50.0
@@
Mulberry Yao et al. (2024) 2B 51.3 77.7 51.7 -
Astar (Ours) 2B 51.7 78.3 49.6 33.7
@@
Insight-V Dong et al. (2024d) 7B 61.5 81.5 59.9 -
AR-MCTS Dong et al. (2024a) 7B - - 64.1 -
Mulberry Yao et al. (2024) 7B 61.3 83.9 63.1 -
LLaVA-CoT OpenAI (2024) 11B 57.6 - 54.8 -
LlamaV-o1 Thawakar et al. (2025) 11B 59.6 - 54.4 -
URSA Luo et al. (2025) 7B - - 59.8 45.7
Virgo Luo et al. (2025) 7B - - - 37.5
Astar (Ours) 7B 61.7 83.9 63.5 54.0
D.3 Integration with SFT
To evaluate the versatility of the AStar framework, we conducted experiments on both pretrained base models and their supervised fine-tuning (SFT) counterparts. Table 13 presents a comprehensive comparison of model performance across different configurations. Our results demonstrate that the AStar framework consistently yields substantial improvements over the zero-shot performance reported in the original papers, regardless of whether it is applied to pretrained or SFT models.
Table 13: Integration With SFT on MathVision. We evaluate the AStar framework using Qwen2-VL-2B and Qwen2-VL-7B as backbone models, comparing both pre-trained base models and SFT variants across different difficulty levels.
Backbone 1 2 3 4 5 Average
Qwen2-VL-2B-Base 20.8 20.7 21.4 17.8 10.2 18.2
Qwen2-VL-2B 25.0 20.0 15.6 16.7 17.6 22.3
Qwen2-VL-7B-Base 26.8 20.6 23.7 26.7 22.4 23.4
Qwen2-VL-7B 26.5 21.3 26.2 29.8 23.8 25.2
[…]
D.4 Weak-to-Strong Generalization

### v3

@@ C.1 Action Space
  • Divide and Conquer ($a_{5}$): Breaking down a complex visual reasoning problem into several smaller subproblems and progressively solving them to achieve the overall solution.
  • Self-Reflection ($a_{6}$): Engages in timely reflection of prior solutions according to text and image sub-regions, and implements necessary refinement during the reasoning process to ensure accuracy.
C.2 Reward Value in MCTS
To simplify the construction of thought cards via the MCTS iteration, we avoid introducing an external reward model to score each step. Given the current skepticism regarding the self-rewarding capabilities of language models, alternative methods are necessary. Inspired by the principle that actions leading to correct answers should be rewarded more frequently, we aim to increase their likelihood of selection in future MCTS tree expansions. Following prior work [66], we define the reward value as the likelihood of self-consistency via majority voting. Note that this principle applies only to leaf nodes. The Q-values of intermediate nodes are initially set to 0 and are dynamically updated during the backpropagation process, as described in Equation 1.
Table 8: The statistics of benchmarks used in this paper. Our evaluation encompasses three tasks: Visual Perception and Understanding (ChartQA, MMStar, MMMU, BLINK), Mathematical Reasoning (MathVista, MathVerse, MathVision), and Commonsense and Scientific Reasoning (GAOKAO-MM).
@@ C.2 Reward Value in MCTS
Visual Perception and Understanding ChartQA [39] chart understanding and reasoning
MMStar [40] 6 core capabilities, like scientific reasoning
MMMU [41] 30 subjects and 183 subfields, like Art and Business
BLINK [42] 14 visual perception tasks, like visual correspondence
Mathematical Reasoning MathVista [36] 12 core capabilities, like arithmetic reasoning
MathVerse [37] 6 distinct vision-text-dominant
MathVision [5] 16 mathematical domains
Commonsense and Scientific Reasoning GAOKAO-MM [38] 8 subjects, like history and Geography
Appendix D Detailed Experimental Setup
D.1 Benchmarks and Datasets
@@
  • ChartQA [39] is a comprehensive benchmark for chart-based question answering that emphasizes visual and logical reasoning capabilities. The dataset comprises 9,608 human-authored questions and 23,111 machine-generated questions derived from human-written chart summaries. It incorporates 20,882 real-world charts from diverse sources including Statista, Pew Research, Our World in Data, and OECD, spanning multiple domains and visualization styles. ChartQA addresses prior dataset limitations through its focus on human-authored questions, real-world charts, and open-vocabulary responses. We utilize this dataset to evaluate our method’s performance on visual chart question-answering tasks commonly encountered in real-world scenarios.
@@
  • MMStar [40] serves as a vision-critical multimodal benchmark for evaluating MLLM capabilities. It addresses two fundamental limitations in existing evaluations: the redundancy of visual inputs and potential training data contamination. Comprising 1,500 meticulously curated challenge samples across six core competencies and 18 evaluation axes, the benchmark undergoes rigorous validation to ensure visual necessity and minimal data leakage. We employ this comprehensive dataset to assess our method’s general reasoning capabilities.
  • MMMU [41] comprises 11,500 meticulously collected multimodal questions from college exams, quizzes, and textbooks, covering six core disciplines: Art &\&& Design, Business, Science, Health &\&& Medicine, Humanities &\&& Social Science, and Tech &\&& Engineering. These questions span 30 subjects and 183 subfields, incorporating 30 highly heterogeneous image types such as charts, diagrams, maps, tables, music sheets, and chemical structures. Unlike existing benchmarks, MMMU focuses on multi-discipline multimodal understanding with domain-specific knowledge, challenging models to perform tasks akin to those faced by experts. Given the interleaved image-text format of the examples and disciplines, these capabilities are considered a kind of Interleaved Image-Text Understanding. We utilize this dataset to assess our method’s performance across a broad spectrum of real-world, expert-level visual perception and understanding.
  • BLINK [42] is a benchmark that evaluates multimodal large language models (MLLMs) on core visual perception abilities not commonly assessed in other evaluations. It reformats 14 classic computer vision tasks into 3,807 multiple-choice questions, each paired with single or multiple images and visual prompts. Tasks include relative depth estimation, visual correspondence, forensics detection, and multi-view reasoning.
@@
  • MathVista [36] is a mathematical visual benchmark consisting of 6,141 examples. These examples are divided into two subsets: testmini with 1,000 annotated examples and test comprising 5,141 problems without public answers. We conduct our evaluation on the testmini subset to assess visual comprehension and compositional reasoning capabilities.
@@
  • MathVerse [37] is a comprehensive and specialized visual mathematics benchmark for assessing multimodal mathematical reasoning capabilities of MLLMs. The benchmark contains 2,612 visual math problems, combining 1,236 newly collected problems from public repositories with 1,376 problems from existing benchmarks. Human annotators have transformed each problem into six distinct versions-text-dominant, text-lite, text-only, vision-intensive, vision-dominant, and vision-only—each offering different levels of multimodal information. To ensure a fair comparison, we conduct evaluations on the testmini set of MathVerse, and construct seed data from the non-overlapping portion of the test set. Note that, MathVerse has no intersection with other tested benchmarks (e.g., GAOKAO-MM), thus preventing data leakage. We utilize this dataset to evaluate AStar’s performance across varying degrees of multimodal integration.
@@
  • MathVision [5] constitutes a meticulously compiled collection of 3,040 mathematics problems, each incorporating visual elements from authentic mathematics competitions. The dataset spans 16 mathematical domains and features a five-tier difficulty classification system. We employ this dataset to evaluate our model’s performance on challenging reasoning tasks.
@@
  • GAOKAO-MM [38] represents a multimodal evaluation framework derived from the Chinese National College Entrance Examination (Gaokao). It encompasses eight academic subjects and includes twelve categories of images, such as diagrams, function graphs, maps, and photographs. The benchmark aims to evaluate models’ abilities to understand and reason over diverse multimodal content, reflecting the complexity and breadth of knowledge. We leverage this dataset to assess AStar’s cross-domain generalization capabilities.
D.2 Comparison Baselines
@@
Open-source general MLLMs: Recent advances include Qwen2-VL [45] and InternVL2 [46], which demonstrate exceptional capabilities in complex reasoning and mathematical problem-solving.
@@
      G-LLaVA [56]: Introduces the Geo170K dataset, comprising over 170K geometric image-caption and question-answer pairs, to enhance MLLMs’ geometric problem-solving capabilities.
@@
      Math-LLaVA [47]: Develops the MathV360K dataset by collecting 40K high-quality images with question-answer pairs and synthesizing an additional 320K pairs, then fine-tunes the LLaVA-1.5 model to improve multimodal mathematical reasoning.
@@
      MultiMath [60]: Constructs the MultiMath-300K dataset, encompassing K-12 level mathematical problems with image captions and step-wise solutions, and trains the MultiMath-7B model to bridge visual and mathematical reasoning.
@@
      Math-PUMA [57]: Proposes a three-stage Progressive Upward Multimodal Alignment methodology to enhance MLLMs’ mathematical reasoning skills, focusing on aligning visual and textual modalities through a structured training process.
@@
      LlamaV-o1 [27]: Introduces a multimodal visual reasoning model trained using a multi-step curriculum learning approach, organizing tasks progressively to facilitate incremental skill acquisition and problem-solving.
@@
      LLaVA-CoT [10]: Develops a vision-language model designed to conduct autonomous multistage reasoning, employing a structured approach that includes summarization, visual interpretation, logical reasoning, and conclusion generation.
@@
      URSA [13]: Proposes a three-module synthesis strategy integrating CoT distillation, trajectory-format rewriting, and format unification, resulting in the MMathCoT-1M dataset and the URSA-7B model, which demonstrates state-of-the-art performance on multiple multimodal mathematical benchmarks.
    3. c.
      Virgo [12]: Explores a straightforward approach to implementing multimodal slow-thinking systems by fine-tuning a capable MLLM with a small amount of textual long-form thought data, demonstrating that such reasoning processes can be effectively transferred to MLLMs.
@@
      InfiMM-Math [59]: Introduces InfiMM-WebMath-40B, a high-quality dataset comprising 24 million web pages, 85 million associated image URLs, and 40 billion text tokens, aiming to improve MLLMs’ mathematical reasoning through large-scale multimodal pre-training.
@@
    Reinforcement Learning with Verifiable Rewards:
@@
      LMM-R1 [20]: Proposes a two-stage rule-based reinforcement learning framework, LMM-R1, which enhances the reasoning capabilities of 3B-parameter Large Multimodal Models (LMMs) by first strengthening foundational reasoning using text-only data, then generalizing these abilities to multimodal domains. It achieves significant performance improvements even with limited data.
    2. b.
      MM-Eureka [48]: Develops the multimodal reasoning model MM-Eureka, which successfully extends large-scale rule-based reinforcement learning to multimodal settings. It reproduces key characteristics of DeepSeek-R1, such as steady increases in accuracy reward and response length during training, and demonstrates superior data efficiency across multiple benchmarks.
    3. c.
      R1-VL [18]: Introduces Step-wise Group Relative Policy Optimization (StepGRPO), a novel online reinforcement learning framework that enhances MLLMs’ reasoning abilities through dense step-wise rewards. By incorporating Step-wise Reasoning Accuracy Reward (StepRAR) and Step-wise Reasoning Validity Reward (StepRVR), R1-VL achieves superior performance on various benchmarks by encouraging more structured and logically consistent reasoning processes.
Advanced closed-source MLLMs: Leading proprietary models including GPT-4V [49], GPT-4o [2], and Gemini-1.5-Pro [73] demonstrate exceptional capabilities in multimodal understanding and task-solving, setting benchmarks for open-source alternatives.
Multimodal tree-based methods: Recent works incorporate explicit search mechanisms into multimodal reasoning. AR-MCTS [14] enhances reasoning by combining Monte Carlo Tree Search (MCTS) with active retrieval, improving reasoning space diversity and reliability. Mulberry [15] leverages multi-model collaboration through MCTS’s four iterative operations (selection, expansion, simulation, backpropagation) to identify optimal reasoning paths.
D.3 Implementation Details
@@
In the thought card construction stage, , during multimodal input-based MCTS iteration, we implement an early termination strategy based on self-consistency for enhanced efficiency. Building on the observation that repeated action sampling at the same state often indicates successful task completion, we terminate simulation early when the model’s consistency score exceeds a threshold $c$ (i.e., $SC(s)>c$). Our experiments show that using a parallel strategy on an A100-80G, acquiring reasoning paths for 500 samples takes only approximately 50 minutes, making this computational overhead minimal compared to large-scale training approaches.
Additionally, we use Problem Condition Complexity (PCC) and Text-Image Semantic (TIS) for problem attributes clustering in card construction; these computational overheads are also manageable. Specifically, this overhead consists of two components:
  • PCC Cost: The overhead comes from using an MLLM to derive prior known conditions. A lightweight Qwen2-VL-2B model is sufficient for this process, keeping costs manageable.
  • TIS Cost: The CLIP-ViT-B-16 (149M) model [34] is computationally efficient. Furthermore, semantic embeddings for 500 seed samples can be pre-computed once (averaging less than 1s per sample), which is a fixed computational overhead. As test samples increase, the amortized CLIP encoding cost per sample diminishes, keeping the actual overhead manageable.
@@
Appendix E Supplementary Results
E.1 Detailed Results on Multimodal Reasoning Benchmarks
@@
LLaVA-1.5-13B [53] 13B 25.7 19.6 28.6 17.6 10.8 27.8 33.6 22.9
InternLM-XComposer2-VL-7B [54] 7B 47.8 32.0 51.6 30.5 13.5 43.8 37.7 62.8
SPHINX-MoE [52] 8$\times$ 7B 42.3 31.7 41.6 30.5 16.2 27.1 50.8 50.8
DeepSeek-VL [55] 7B 34.9 29.2 38.8 27.2 18.9 43.1 35.3 33.2
InternVL2-8B [46] 8B 58.3 59.8 56.4 60.3 10.8 30.6 59.0 68.8
Qwen2-VL [45] 7B 58.9 44.1 57.5 43.1 24.3 41.7 66.4 75.1
@@
G-LLaVA [56] 7B 25.1 36.0 19.4 37.6 15.2 17.7 21.0 15.1
Math-LLaVA [47] 13B 46.6 51.5 40.7 56.2 23.3 34.7 47.7 42.3
Multimath-7B [60] 7B 50.0 61.9 42.2 64.9 23.3 32.6 42.6 49.2
Math-PUMA-Qwen2-7B [57] 7B 47.9 47.7 46.2 47.3 21.6 32.6 42.6 55.8
URSA-7B 7B 59.8 74.0 53.5 77.4 21.6 35.4 58.2 57.1
AStar (Ours, Training-free Reasoning) 7B 64.2 69.0 63.8 71.7 59.5 60.0 48.2 69.1
@@
E.2 Comparison with Strong Baselines
@@
We also benchmark our approach against recent powerful multimodal reasoning methods, including LLaVA-CoT [10] and Virgo [12]. As shown in Table 12, comparative experiments across different model scales demonstrate that our AStar reasoning framework, which effectively integrates MLLMs’ inherent reasoning capabilities with external reasoning guidelines, achieves superior performance with minimal prior data. On challenging datasets like MathVerse, our method exhibits substantial improvements, with the 2B model achieving comparable performance to the extensively trained Virgo 7B model. Further analysis reveals two key insights: First, methods relying on vast solution spaces often struggle to identify appropriate reasoning paths, analogous to finding a needle in a haystack. Second, approaches dependent on large-scale training data typically face difficulties in fully capturing complex long-chain reasoning patterns through implicit learning. In contrast, our method adaptively identifies suitable reasoning strategies based on problem complexity, enabling efficient inference across diverse scenarios.
[…]
@@
GPT-4o-0513 [2] 63.8 50.2 30.4 48.1
GPT-4V [49] 49.9 54.4 24.8 43.1
Gemini-1.5-Pro [73] 63.9 35.3 19.2 39.5
Claude-3.5-Sonnet [75] 67.7 - - -
@@
Qwen2-VL-72B [45] 70.5 41.3 25.9 45.9
InternVL2.5-26B [46] 67.7 40.1 28.0 45.3
InternVL2-Llama3-76B [46] 65.5 42.8 23.7 44.0
Astar (Qwen2.5-7B) 64.2 53.9 32.7 50.3
Astar (Qwen2-VL-7B) 61.7 47.5 26.8 45.3
@@
Mulberry [15] 2B 51.3 77.7 51.7 -
Astar (Qwen2-VL-2B) 2B 52.0 78.6 49.6 35.6
@@
Insight-V [11] 7B 61.5 81.5 59.9 -
AR-MCTS [14] 7B - - 64.1 -
Mulberry [15] 7B 61.3 83.9 63.1 -
LLaVA-CoT [2] 11B 57.6 - 54.8 -
LlamaV-o1 [27] 11B 59.6 - 54.4 -
URSA [13] 7B 55.4 - 59.8 45.7
Virgo [13] 7B - - - 37.5
Astar (Qwen2.5-7B) 7B 62.3 84.2 64.2 53.9
Astar (Qwen2-VL-7B) 7B 61.7 83.9 61.7 47.5
E.3 Weak-to-Strong Generalization